OpenStudy (anonymous):

Plug following polynomials into the quadratic equation. Show all steps: a) y=0.1269x^2-1.5741x+6.792 b) y=0.0184x^2+1.0496x-7.4804

OpenStudy (anonymous):

you gotta be kidding

OpenStudy (anonymous):

I kid you, not.

OpenStudy (anonymous):

what about plugging it into the equation dont you understand?

OpenStudy (anonymous):

\[\frac{1.5741\pm\sqrt{(1.5741)^2-4\times .1269\times 6.792}}{2\times.1269}\]

OpenStudy (anonymous):

is the first one.

OpenStudy (anonymous):

I understand it....I just don't have time to do it cause I'm doing something else at work. Steps, please?